# A Sober Look at Progress in Language Model Reasoning: Pitfalls and Paths to Reproducibility

### v1

@@ 3.3 Variance from Hardware and Software Factors
Evaluation across different Python frameworks. Evaluation results can vary based on the framework used, due to differences in prompt templates, inference engines (e.g., vLLM (Kwon et al., 2023)), and response extraction strategies (e.g., MathVerify). For example: lighteval is used by OpenRS (Dang & Ngo, 2025), evalchemy (Guha et al., 2024) is used by models like OpenThinker and Bespoke-Stratos, other frameworks include lm-evaluation-harness (Gao et al., 2024b) and evalscope (Alibaba ModelScope Community, ).
To assess this impact, we compare lighteval and evalchemy, keeping all other variables fixed: model, dataset, hardware, decoding parameters, and random seeds (3 per model). For a fair comparison, we evaluated two models, DeepSeek-R1-Distill-1.5B and S1.1-7B, at their default temperature and top_p parameter values on a single GPU. We present results averaged over three seeds for higher robustness. As shown in Table 2, framework-induced differences are generally small (1–2pp) but can still affect model rankings in tightly clustered scenarios.
@@
3.4 Effect of Prompt Format and Context Length
Prompt Format. Prompt formatting has a measurable impact on accuracy. As shown in Figure 10, models perform best when using math-specific prompts and their native chat templates. Omitting templates leads to performance drops, particularly for instruction-tuned models. We compare accuracy under three different prompt settings (see Table 5): (1) a math-specific prompt formatted using the model’s chat template, (2) only the model’s chat template with no additional prompt, and (3) no template at all, i.e., the question without any special tokens or instructions. Interestingly, while base models like Qwen2.5-Math may benefit from prompt-free setups (Liu et al., 2025b), instruction-tuned models rely heavily on format alignment. Thus, maintaining consistent and format-aware prompting is essential for maximizing instruction-tuned model performance.
@@ 4.1 Recommendations: Which practices to adopt?
Variance Estimates: For small benchmarks (e.g., AIME’24), run evaluations with at least ten random seeds. Report the mean and standard deviation to quantify uncertainty and assess the statistical significance of performance differences.
@@
Context Length and Prompt Template Selection: Ensure the context length is sufficiently large—especially for models with long reasoning chains—to avoid premature truncation and under-reported accuracy. For instruction-tuned models, always use the appropriate chat template to match the expected input format.
Robust Answer Matching: We strongly recommend using a resilient answer extraction pipeline that handles parsing issues and evaluates expression equivalence, rather than relying on exact string matching. This reduces the likelihood of spurious gains from formatting artifacts.
@@ 4.2 Standardization Procedure
We adopt a largely consistent experimental setup with prior work, with the key difference being our use of publicly accessible cloud instances from Runpod222https://www.runpod.io/pricing. Each instance is equipped with a single A100 PCIe GPU, 8 vCPUs, and 128 GB of RAM. We evaluate all models listed in Table 3 across six benchmarks: AIME’24 (AI-MO, ), AIME’25 (Lin, 2025), AMC’23 (Knovel Engineering, 2025), MATH500 (HuggingFaceH4, 2024), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). All experiments are conducted using the LightEval framework (Fourrier et al., 2023) (0.8.1) with a vLLM backend, repeated across ten random seeds for AIME’24, AIME’25, AMC’23 and three random seeds for the rest. Depending on the base model architecture, we set the maximum number of new tokens (e.g., 4096 for QwenMath-based models), apply optimal hyperparameters, and use the appropriate chat template. LightEval’s LaTeX-based answer extraction and evaluation pipeline ensures reliable and consistent result parsing and correctness matching, similar to math-verify.
@@ 4.3 A Sober Look: Results
II-Thought (Intelligent Internet, 2025) | 32.05.9 | 24.04.1 | 79.55.1 | 86.60.6 | 31.70.6 | 54.90.4 |
FastCuRL (Song et al., 2025) | 36.34.3 | 27.03.7 | 78.84.1 | 87.91.2 | 30.81.4 | 56.50.6 |
DeepScaleR (Luo et al., 2025) | 37.06.6 | 30.34.3 | 76.24.6 | 87.81.0 | 31.01.5 | 55.51.1 |
@@
Light-R1 (Wen et al., 2025a) | 53.04.8 | 41.03.5 | 90.03.1 | 93.50.5 | 41.31.3 | 68.01.2 |
@@
Math (Base) (Yang et al., 2024b) | 11.33.6 | 5.72.7 | 44.04.9 | 51.75.5 | 11.32.2 | 26.00.6 |
Oat-Zero (Liu et al., 2025a) | 16.03.2 | 6.73.4 | 52.52.9 | 73.51.7 | 26.30.8 | 37.21.3 |
Math (Instruct) (Yang et al., 2024b) | 12.01.7 | 11.75.7 | 54.85.3 | 74.70.5 | 26.71.8 | 37.90.2 |
@@
Math (Base) (Yang et al., 2024b) | 20.73.8 | 8.73.9 | 56.25.7 | 64.30.5 | 17.31.9 | 29.00.5 |
SimpleRL-Zoo (Zeng et al., 2025b) | 22.75.2 | 10.73.4 | 62.23.6 | 76.91.8 | 30.12.8 | 39.30.6 |
LIMR (Li et al., 2025a) | 30.73.2 | 7.83.3 | 62.23.4 | 76.50.4 | 34.91.3 | 39.30.9 |
Oat-Zero (Liu et al., 2025a) | 28.03.1 | 8.82.5 | 66.23.6 | 79.40.3 | 34.41.4 | 43.81.1 |
Math (Instruct) (Yang et al., 2024b) | 15.73.9 | 10.73.8 | 67.03.9 | 82.90.1 | 35.00.6 | 41.30.9 |
@@
Qwen (Base) (Yang et al., 2024a) | 0.00.0 | 0.00.0 | 2.52.5 | 3.31.5 | 1.80.4 | 1.50.5 |
SimpleRL-Zoo (Zeng et al., 2025b) | 0.31.1 | 0.31.1 | 13.24.7 | 12.06.5 | 4.02.4 | 4.22.0 |
Qwen (Instruct) (Yang et al., 2024a) | 1.31.7 | 0.71.4 | 26.24.8 | 57.51.1 | 19.41.3 | 20.31.1 |
@@
Qwen (Base) (Yang et al., 2024a) | 3.33.3 | 0.00.0 | 30.09.0 | 64.61.0 | 25.70.9 | 30.11.2 |
SimpleRL-Zoo (Zeng et al., 2025b) | 14.02.1 | 4.32.7 | 58.01.6 | 77.90.8 | 33.00.2 | 39.00.1 |
Open Reasoner Zero (Hu et al., 2025) | 19.72.9 | 15.72.7 | 59.54.5 | 83.91.1 | 31.61.3 | 47.61.7 |
Qwen (Instruct) | 12.33.2 | 7.33.4 | 52.84.8 | 77.11.2 | 34.91.0 | 38.71.0 |
Based on: Qwen2.5 7B (SFT) | ||||||
Qwen (Instruct) (Yang et al., 2024a) | 12.33.2 | 7.33.4 | 52.84.8 | 77.11.2 | 34.91.0 | 38.71.0 |
Eurus2 Prime (Cui et al., 2025) | 17.82.2 | 14.01.7 | 63.03.9 | 80.10.1 | 37.51.0 | 43.90.3 |
s1.1 (Muennighoff et al., 2025) | 19.03.2 | 21.05.5 | 59.53.7 | 80.80.6 | 37.51.1 | 48.21.4 |
Bespoke Stratos (Bespoke Labs, 2024) | 20.34.3 | 18.04.8 | 60.24.9 | 84.70.5 | 39.11.3 | 51.91.1 |
OpenThinker (Team, 2025) | 30.56.2 | 26.04.4 | 71.43.9 | 88.31.4 | 37.93.8 | 55.61.4 |
OpenR1 (Face, 2025) | 48.38.9 | 35.54.2 | 86.04.5 | – | – | – |
OpenThinker2 (Team, 2025) | 53.04.6 | 41.05.0 | 87.03.5 | 81.60.7 | 33.90.2 | 46.91.3 |
@@
RL-training on R1-Distill We evaluated several reinforcement learning (RL) approaches (e.g., GRPO) using the DeepSeek R1-Distill-1.5B model. We first observe that none of the L1 models (Aggarwal & Welleck, 2025) outperformed the original DeepSeek R1-Distill baseline — an expected outcome given that L1 training prioritized smaller output length over accuracy. OpenRS (Dang & Ngo, 2025) reported strong gains (10–15%) on AIME, AMC, and OlympiadBench. However, our replication showed no statistically significant improvements over the R1 - Distill baseline. Same case held for Still-3 and Light-R1 model, which showed no significant improvement over the R1-Distill baseline. II-Thought and FastCurl yield modest improvements across benchmarks, especially over AIME’24 but the observed gains did not carry over significantly to AIME’25 indicating overfitting to existing benchmarks. Only DeepscaleR demonstrated robust, significant improvements across benchmarks.
@@
RL Training on Qwen2.5 Math and Base Models: We next analyze RL training applied to the Qwen2.5 Base and Qwen2.5 Math Base models, a trend trying to replicate gains by Deepseek-R1 Zero. Unlike the R1-Distill results, RL training with Oat-Zero, LIMR, and SimpleRL-Zoo consistently produced statistically significant gains over the base model, especially across Math500, Minerva and OlympiadBench benchmarks. This indicates that RL-based approaches can indeed offer substantial improvements given a base model instead of a distilled R1 model. However, these gains remained smaller than those achieved via instruction tuning in the original Qwen papers, suggesting that instruction tuning alone may be sufficient to far surpass current gains from RL methods in this setting. We also observed that the improvements on AIME’24 were also significant, but did not carry over to AIME’25 indicating a troubling overfitting trend. Notably, Open Reasoner-Zero-7B was the only RL-trained model to consistently outperform the instruct-tuned baseline by large margins across all benchmarks.
@@
Effectiveness of Supervised Finetuning. We assessed supervised finetuning methods like s1.1, Eurus2 Prime, Bespoke Stratos, OpenR1 and OpenThinker models, which further refine instruction-tuned models using reasoning traces. Supervised methods consistently outperformed the instruct-tuned baseline across all benchmarks (even Minerva) and generalized comparatively well to AIME’25. The performance improvements from OpenThinker were especially notable. These results underscore the maturity and effectiveness of SFT when training recipes are scaled to large datasets.
@@
Overfitting and Generalization We now examine the overfitting by comparing performance on AIME’24 versus the more challenging AIME’25. RL-trained models showed a pronounced performance drop between the two, indicating overfitting to the training distribution. In contrast, supervised fine-tuning (SFT) models maintained consistent improvements, suggesting better generalization. Openthinker2 showed significant degradation compared to Openthinker across benchmarks not provided in their blogpost, indicating overfitting via data-curation. This highlights a gap in current evaluation protocols, and a need to assess out-of-distribution generalization for reasoning models.
4.4 Do Discovered Phenomena Replicate? A Detailed Analysis.
@@
We further investigate two recently noted phenomena to see if they replicate in our experiments: (1) how response length correlates with performance, and (2) the decline in response diversity following reasoning-focused training.
4.4.1 Are Incorrect Responses Longer?

### v2

@@ 3.4 Variance from Hardware & Software Factors: A Critical Consideration
Mitigation strategies. We conducted extensive experiments to isolate and quantify hardware-induced variability, even when using identical GPU types and software stacks. Our investigation had 3 stages, each attempting to eliminate additional sources of variation:
Stage 1: Standardized Docker across different A100 clusters. We first evaluated models using our standardized Docker container on A100 GPUs across two platforms: our internal cluster and Runpod (a cloud GPU service). Despite identical hardware specifications and containerized environments, we observed notable performance differences. For instance, Qwen2.5-Math-1.5B achieved 7.3%3.8 on AIME’24 on our cluster versus 11.3%3.6 on Runpod; a 4 percentage point gap that exceeds the standard deviation.
Stage 2: Enforcing CUDA determinism. Suspecting non-deterministic GPU operations, we enforced strict determinism by setting torch.use_deterministic_algorithms(True), fixing CUDA seeds, and disabling cuDNN benchmarking. Comparing A100 and H100 clusters, variance decreased but persisted: OpenThinker2-7B scored 53.0%4.6 on A100 versus 57.1%5.2 on H100 for AIME’24. Even with deterministic algorithms, hardware architectural differences and vLLM’s backend variations (vLLM Contributors, 2024) continued to introduce discrepancies.
Stage 3: Identical hardware, updated stack. Finally, we updated all components (vLLM, LightEval, math-verify) and ran multiple evaluations on identical hardware. Remarkably, even consecutive runs on the same A100 cluster produced variations: Bespoke-Stratos-7B ranged from 19.3% to 23.0% on AIME’24 across runs. These persistent differences, despite controlling for hardware, software versions, and random seeds highlight deep-seated sources of non-determinism in modern inference stacks, potentially arising from dynamic kernel selection, memory allocation patterns, or floating-point accumulation order (Atil et al., 2025).
As shown in Figure 10, these effects are not merely theoretical but measurably impact benchmark accuracy even when averaging over multiple seeds. This underscores that reproducibility challenges extend beyond simple seed variance to fundamental architectural and systems-level factors (Atil et al., 2025; Yao et al., 2025; He & Lab, 2025).
Evaluation across different Python frameworks. Evaluation results can vary based on the framework used, due to differences in prompt templates, inference engines (e.g., vLLM (Kwon et al., 2023)), and response extraction strategies (e.g., MathVerify). For example: lighteval is used by OpenRS (Dang & Ngo, 2025), evalchemy (Raoof et al., 2025) is used by models like OpenThinker and Bespoke-Stratos, other frameworks include lm-evaluation-harness (Gao et al., 2024b) and evalscope (Alibaba ModelScope Community, 2025). Some prior works have also studied the performance differences induced by using different evaluation frameworks for evaluating base language models (Li et al., 2024).
To assess this impact, we compare lighteval and evalchemy, fixing all other variables: model, dataset, hardware, decoding parameters, and random seeds (3 per model). For a fair comparison, we evaluated two models, DeepSeek-R1-Distill-1.5B and S1.1-7B, using default temperature and top_p parameter values on a single GPU. We present results averaged over three seeds for higher robustness. As shown in Table 2, framework-induced differences are generally small (1–2pp) but can still affect model rankings in tightly clustered scenarios.
@@ 4.1 Recommendations: Which practices to adopt?
Variance Estimates: For small benchmarks (e.g., AIME’24), run evaluations with at least 30 random seeds. Report the mean and standard deviation to quantify uncertainty and assess the statistical significance of performance differences.
@@
Context Length and Prompt Template Selection: Ensure the context length is sufficiently large, especially for models with long reasoning chains to avoid premature truncation and under-reported accuracy. For instruction-tuned models, always use the appropriate chat template to match the expected input format.
Robust Answer Extraction: We strongly recommend using a resilient answer extraction pipeline that handles parsing issues and evaluates expression equivalence, rather than relying on exact string matching. This reduces the likelihood of spurious gains from formatting artifacts.
@@ 4.2 Standardization Procedure
We adopt a largely consistent experimental setup with prior work, with the key difference being our use of publicly accessible cloud instances from Runpod222https://www.runpod.io/pricing. Each instance is equipped with a single A100 PCIe GPU, 8 vCPUs, and 128 GB of RAM. We evaluate all models listed in Table 3 across six benchmarks: AIME’24 (AI-MO, 2024a), AIME’25 (Lin, 2025), AMC’23 (Knovel Engineering, 2025), MATH500 (HuggingFaceH4, 2024), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). All experiments are conducted using the LightEval framework (Fourrier et al., 2023) (0.8.1) with a vLLM backend, repeated across ten random seeds for AIME’24, AIME’25, AMC’23 and three random seeds for the rest. While our analysis in section 3.2.1 identifies 30 seeds as optimal for variance stabilization, we use 10 seeds for AIME/AMC benchmarks as a practical compromise between rigor and computational cost. This still represents a significant improvement over the single-seed evaluations common in prior work, and we report standard deviations to quantify remaining uncertainty. Depending on the base model architecture, we set the maximum number of new tokens (e.g., 4096 for QwenMath-based models), apply optimal hyperparameters, and use the standardized chat template except for base models. LightEval’s LaTeX-based answer extraction and evaluation pipeline ensures reliable and consistent result parsing and correctness matching, similar to math-verify.
Statistical Significance Testing. To rigorously assess whether RL-trained models genuinely outperform their baselines, we employ paired sample-wise statistical tests rather than simply comparing mean accuracies. For each model-baseline pair, we: (1) compute the average accuracy per problem instance across all random seeds, (2) align these sample-wise accuracies between the RL/SFT model and the baseline from which it was trained, and (3) conduct one-sided paired tests to evaluate whether the RL model achieves systematically higher performance. We employ both parametric (paired t-test) and non-parametric (Wilcoxon signed-rank) tests to ensure robustness to distributional assumptions. The null hypothesis is that there is no improvement from RL/SFT training, with the alternative hypothesis that the RL/SFT model performs better than its baseline. We report results as statistically significant at two thresholds: (marked with ∗) and (marked with ∗∗). This conservative approach ensures that reported improvements represent genuine performance gains rather than statistical artifacts from multiple testing or random variation.
@@ 4.3 A Sober Look: Results
| ZR1-1.5B (Zyphra, 2025) | 30.34.6 | 24.03.8 | 78.84.0 | 86.60.9 | 32.10.9 | 56.40.4∗∗ |
| II-Thought (Intelligent Internet, 2025) | 32.05.9 | 24.04.1 | 79.55.1∗ | 86.60.6 | 31.70.6 | 54.90.4∗ |
| DeepScaleR (Luo et al., 2025) | 37.06.6 | 30.34.3∗∗ | 76.24.6 | 87.81.0∗∗ | 31.01.5 | 55.51.1∗∗ |
| FastCuRL (Song et al., 2025) | 36.34.6∗ | 26.53.7 | 76.94.8 | 87.51.0∗∗ | 30.81.6 | 56.70.8∗∗ |
| DisCO-logL (Li et al., 2025a) | 39.04.5 | 31.74.8∗ | 78.83.4∗ | 87.41.1∗ | 32.71.3 | 58.60.3∗∗ |
| Nemotron-RR (Liu et al., 2025a) | 48.36.3∗∗ | 33.05.1∗ | 87.04.8∗∗ | 91.00.5∗∗ | 37.51.6∗∗ | 62.80.9∗∗ |
@@
| Sky-T1 (mini) (Team, 2025a) | 51.34.5 | 39.05.7 | 90.03.3 | 93.31.0 | 41.81.2 | 67.70.9 |
| DisCO-logL (Li et al., 2025a) | 51.33.9 | 41.34.5 | 92.22.2 | 94.01.3 | 44.71.1∗∗ | 67.60.3 |
| Skywork-OR1 (He et al., 2025) | 60.74.7∗ | 49.76.2∗ | 94.03.8 | 95.70.9∗ | 43.01.0∗ | 73.60.9 ∗∗ |
| Based on: Deepseek R1 Distill Qwen 7B (SFT) | ||||||
| Light-R1 (Wen et al., 2025) | 53.04.8 | 41.03.5 | 90.03.1 | 93.50.5 | 41.31.3 | 68.01.2 |
@@
| Math (Base) (Yang et al., 2024) | 11.33.6 | 5.72.7 | 44.04.9 | 51.75.5 | 11.32.2 | 26.00.6 |
| Oat-Zero (Liu et al., 2025b) | 16.03.2 | 6.73.4 | 52.52.9 | 73.51.7∗∗ | 26.30.8∗∗ | 37.21.3∗∗ |
| Math (Instruct) (Yang et al., 2024) | 12.01.7 | 11.75.7∗ | 54.85.3 | 74.70.5∗∗ | 26.71.8∗∗ | 37.90.2∗∗ |
| LUFFY-Zero (Yan et al., 2025a) | 15.94.9 | 11.72.8 | 54.55.4 | 77.61.2∗∗ | 27.20.4∗∗ | 42.30.5∗∗ |
@@
| Math (Base) (Yang et al., 2024) | 20.73.8 | 8.73.9 | 56.25.7 | 64.30.5 | 17.31.9 | 29.00.5 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 22.75.2 | 10.73.4 | 62.23.6 | 76.91.8∗∗ | 30.12.8∗∗ | 39.30.6∗∗ |
| LIMR (Li et al., 2025b) | 30.73.2 | 7.83.3 | 62.23.4 | 76.50.4∗∗ | 34.91.3∗∗ | 39.30.9∗∗ |
| Oat-Zero (Liu et al., 2025b) | 28.03.1 | 8.82.5 | 66.23.6 | 79.40.3∗∗ | 34.41.4∗∗ | 43.81.1∗∗ |
| Sky-T1 (Team, 2025a) | 19.32.6 | 21.03.9∗ | 67.23.6∗∗ | 85.20.5∗∗ | 34.71.1∗∗ | 52.10.8∗∗ |
| Math (Instruct) (Yang et al., 2024) | 15.73.9 | 10.73.8 | 67.03.9 | 82.90.1∗∗ | 35.00.6∗∗ | 41.30.9∗∗ |
| LUFFY-Zero (Yan et al., 2025a) | 26.76.5 | 23.35.7∗ | 73.53.9∗∗ | 87.90.1∗∗ | 34.12.2∗∗ | 54.90.3∗∗ |
@@
| Qwen (Base) (Qwen et al., 2025) | 0.00.0 | 0.00.0 | 2.52.5 | 3.31.5 | 1.80.4 | 1.50.5 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 0.31.1 | 0.31.1 | 13.24.7∗ | 12.06.5∗∗ | 4.02.4∗ | 4.22.0∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 1.31.7 | 0.71.4 | 26.24.8∗∗ | 58.11.4∗∗ | 19.41.3∗∗ | 20.40.5∗∗ |
@@
| Qwen (Base) (Qwen et al., 2025) | 8.03.2 | 3.73.7 | 35.84.9 | 59.70.3 | 21.41.9 | 27.01.0 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 14.02.1 | 4.32.7 | 58.01.6∗∗ | 77.90.8∗∗ | 33.00.2∗∗ | 39.00.1∗∗ |
| Open Reasoner Zero (Hu et al., 2025) | 19.72.9 | 15.72.7 | 59.54.5∗∗ | 83.91.1∗∗ | 31.61.3∗∗ | 47.61.7∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 12.33.2 | 7.33.4 | 52.84.8∗∗ | 77.30.8∗∗ | 34.91.0∗∗ | 38.90.9∗∗ |
| Based on: Qwen2.5 7B Instruct (SFT) | ||||||
| Qwen (Instruct) (Qwen et al., 2025) | 12.33.2 | 7.33.4 | 52.84.8 | 77.30.8 | 34.91.0 | 38.90.9 |
| Eurus2 Prime (Cui et al., 2025) | 18.71.7 | 14.31.6 | 64.84.0 | 80.10.1∗∗ | 37.51.0∗∗ | 43.90.3∗∗ |
| s1.1 (Muennighoff et al., 2025) | 19.03.2 | 21.05.5∗ | 59.53.7 | 80.80.6∗ | 37.51.1 | 48.21.4∗∗ |
| Bespoke Stratos (Bespoke Labs, 2024) | 20.34.3 | 18.04.8 | 60.24.9 | 84.70.5∗∗ | 39.11.3∗ | 51.91.1∗∗ |
| OpenThinker (Open Thoughts, 2025) | 30.56.2∗∗ | 26.04.4∗∗ | 71.43.9∗∗ | 88.31.4∗∗ | 37.93.8∗∗ | 55.61.4∗∗ |
| OpenR1 (Hugging Face, 2025) | 48.38.9∗∗ | 35.54.2∗∗ | 86.04.5∗∗ | 93.50.7∗∗ | 41.21.3∗∗ | 67.41.3∗∗ |
| OpenThinker2 (Open Thoughts, 2025) | 53.04.6∗∗ | 41.05.0∗∗ | 87.03.5∗∗ | 94.40.7∗∗ | 42.01.5∗∗ | 70.61.0∗∗ |
| OpenThinker3 (Open Thoughts, 2025) | 66.75.2∗∗ | 57.06.4∗∗ | 91.82.6∗∗ | 95.90.6∗∗ | 42.80.8∗∗ | 74.80.8∗∗ |
| Model | AIME’24 | AIME’25 | AMC’23 | MATH500 | Minerva | Olympiad |
| Based on: Qwen2.5 32B (RL) | ||||||
| Qwen (Base) (Qwen et al., 2025) | 8.05.5 | 2.32.2 | 36.56.8 | 62.83.5 | 28.92.2 | 29.52.0 |
| DAPO (Yu et al., 2025) | 43.04.0∗∗ | 32.36.1∗∗ | 88.53.6∗∗ | 91.00.3∗∗ | 40.00.6∗∗ | 60.91.1∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 15.74.7 | 13.04.6 | 60.85.7 | 81.60.5 | 40.31.3 | 46.71.3 |
| Based on: Qwen QwQ 32B (RL) | ||||||
| QwQ-32B (Team, 2025b) | 76.33.3 | 69.04.5 | 96.22.4 | 97.50.6 | 49.00.2 | 78.11.0 |
| INTELLECT-2 (Prime Intellect, 2025) | 75.22.4 | 66.36.6 | 96.02.7 | 97.00.3 | 49.80.2 | 78.00.8 |
| Based on: DeepSeek R1 Distill Qwen 32B (SFT) | ||||||
| Distill R1 (DeepSeek-AI, 2025) | 67.01.9 | 55.35.7 | 96.82.1 | 95.10.7 | 45.11.7 | 73.80.5 |
| TinyR1-Preview (Sun et al., 2025) | 73.34.4 | 66.06.6∗ | 97.82.2 | 97.20.4∗∗ | 48.50.5∗ | 76.31.2∗ |
| Light-R1 DS (Wen et al., 2025) | 76.74.7∗ | 68.75.7∗∗ | 97.01.6 | 97.30.1∗∗ | 48.21.1∗ | 76.90.8∗∗ |
| Based on: Qwen2.5 32B Instruct (SFT) | ||||||
| Qwen (Instruct) (Qwen et al., 2025) | 15.74.7 | 13.04.6 | 60.85.7 | 81.60.5 | 40.31.3 | 46.71.3 |
| Sky-T1-Preview (Team, 2025a) | 37.08.1∗∗ | 22.73.1 | 75.83.5∗∗ | 88.60.5∗∗ | 39.81.2 | 56.30.8∗∗ |
| Bespoke-Stratos (Bespoke Labs, 2024) | 37.75.9∗∗ | 26.37.3∗ | 80.73.9∗∗ | 90.30.3∗∗ | 43.31.5 | 61.51.7∗∗ |
| LIMO (Ye et al., 2025) | 57.84.7∗∗ | 46.35.5∗∗ | 93.02.8∗∗ | 94.61.2∗∗ | 45.22.3∗ | 70.50.9∗∗ |
| s1.1-32B (Muennighoff et al., 2025) | 61.36.9∗∗ | 49.06.9∗∗ | 94.23.9∗∗ | 94.80.5∗∗ | 46.20.8∗∗ | 72.40.7∗∗ |
| OpenThinker (Open Thoughts, 2025) | 68.56.7∗∗ | 50.36.2∗∗ | 95.02.9∗∗ | 95.50.3∗∗ | 46.81.2∗∗ | 71.80.5∗∗ |
| OpenThinker2 (Open Thoughts, 2025) | 71.36.5∗∗ | 61.75.3∗∗ | 95.82.9∗∗ | 96.10.4∗∗ | 46.60.6∗∗ | 75.50.6∗∗ |
| Standalone Models (No tools) | ||||||
| GPT-OSS-20B (medium) (OpenAI, 2025b) | 75.75.0 | 72.75.4 | 97.03.3 | 96.50.8 | 45.11.1 | 77.00.7 |
@@
RL-training on R1-Distill. We evaluated several reinforcement learning (RL) approaches (e.g., GRPO) using the DeepSeek R1-Distill-1.5B model. We first observe that none of the L1 models (Aggarwal & Welleck, 2025) outperformed the original DeepSeek R1-Distill baseline — an expected outcome given that L1 training prioritized smaller output length over accuracy. OpenRS (Dang & Ngo, 2025) reported strong gains (10–15%) on AIME, AMC, and OlympiadBench. However, our replication showed no statistically significant improvements over the R1-Distill baseline. Same case held for Still-3 and Light-R1 models, which showed no significant improvement over the R1-Distill baseline. II-Thought yields modest improvements across benchmarks, especially over AIME’24 but the observed gains did not carry over significantly to AIME’25 indicating overfitting to existing benchmarks. Only DeepscaleR and FastCuRL demonstrated statistically significant improvements across many benchmarks. Notably, recent models like Nemotron-RR deliver the first recipes with robust improvements over the DeepSeek R1-Distill baseline.
@@
RL Training on Qwen2.5 Math and Base Models. We next analyze RL training applied to the Qwen2.5 Base and Qwen2.5 Math Base models, a trend trying to replicate gains by Deepseek-R1 Zero. Unlike the R1-Distill results, RL training with Oat-Zero, LIMR, and SimpleRL-Zoo consistently produced statistically significant gains over the base model, especially across Math500, Minerva and OlympiadBench benchmarks. This indicates that RL-based approaches can indeed offer substantial improvements given a base model instead of a distilled R1 model. However, these gains are often roughly comparable and sometimes slightly higher than achieved via instruction tuning in the original Qwen papers, suggesting that instruction tuning with additional math data alone may be sufficient to far achieve the current gains from RL methods in this setting. We also observed that the improvements on AIME’24 were also significant, but did not carry over to AIME’25 indicating a troubling overfitting trend.
@@
Effectiveness of Supervised Finetuning. We assessed supervised finetuning methods like s1.1, Eurus2 Prime, Bespoke Stratos, OpenR1 and OpenThinker models, which further refine instruction-tuned models using reasoning traces. Supervised methods consistently outperformed the instruct-tuned baseline across all benchmarks (even Minerva) and generalized comparatively well to AIME’25. The performance improvements from OpenThinker series and OpenR1 were especially notable, showcasing the promise of data curation for supervised finetuning (Guha et al., 2025). These results underscore the maturity and effectiveness of SFT when training recipes are scaled to large datasets.
Scaling to Larger Parameters. We extend our analysis to 32B scale models in Table 4. Among R1-Distill variants, Light-R1 DS achieves significant performance improvements, demonstrating that supervised finetuning can further improve even strong distilled models. For Qwen2.5-32B base models, SFT approaches like OpenThinker series achieve strong gains (71.3% and 68.5% on AIME’24 respectively) over the instruct model, further emphasising the importance of high-quality data curation. Notably, DAPO - an RL approach, shows improvements over the base model, and showcases strong performance gains even relative to the Qwen Instruct model. The QwQ-32B models demonstrate state-of-the-art performance (76.3% on AIME’24), and INTELLECT-2 shows no meaningful improvement over QwQ-32B despite additional training. These results reinforce our earlier findings: (1) SFT on reasoning traces remains the most reliable approach for performance gains, and (2) scaling to larger models amplifies these benefits, with 32B models achieving substantially higher absolute performance than their smaller counterparts.
@@
Overfitting and Generalization. We now examine the overfitting by comparing performance on AIME’24 versus the more challenging AIME’25. RL-trained models showed a pronounced performance drop between the two, indicating overfitting to the training distribution. In contrast, supervised fine-tuning (SFT) models maintained consistent improvements, suggesting better generalization. Openthinker2 showed significant degradation compared to Openthinker across benchmarks not provided in their blogpost, indicating overfitting via data-curation. This highlights a gap in current evaluation protocols, and a need to assess out-of-distribution generalization for reasoning models.
Having established these standardized benchmarking results, we now examine whether several recently reported phenomena in reasoning models replicate under our rigorous evaluation framework.
5 Do Discovered Phenomena Replicate? A Detailed Analysis.
@@
We further investigate three recently noted phenomena to see if they replicate in our experiments: (1) how response length correlates with performance, (2) the decline in response diversity following reasoning-focused training and (3) the effect of spurious prompts on performance
5.1 Are Incorrect Responses Longer?